OpenStudy (anonymous):

Use synthetic substitution to evaluate the polynomial for the given value x = -3. P(x) = 2x2 + 7x - 1 P(-3) = _____ Enter a number only.

OpenStudy (anonymous):

is the answer -4?

OpenStudy (anonymous):

2(-3)^2 + 7(-3) -1 = 2(9) - 21 -1 = 18 - 22 = -4 ( check )

OpenStudy (raden):

yes, that's right